Let X and Y be Jointly Normal random variables with: E[Y] = 0, O1, y 2andE[X|Y = y 1 4

Find the joint PDF of X and Y
Answer #1

R D X, Y are jointly Normal Distri Random Variable. EX/Y = y) = 7+ 1., ie. the regression ean equation of x on Y is given byAgain, E (x ly = y) = 4 + 1 > E[ExlY=y)] = E(+1) = EU) + 1 = 1. - E(X) = 1. P = 0.25. E(Y) = 0, y = Jy = 1 So, the joint PDF